OpenStudy (jordanloveangel):

What is the equation of the graphed line? Hint: Determine the slope of the line. y = ? x

OpenStudy (jordanloveangel):

@iishineontheinside

OpenStudy (camerondoherty):

So Slope is \[\frac{ Rise }{ Run }\] So tell me, what is the rise? the run?

OpenStudy (jordanloveangel):

1,-5?

OpenStudy (anonymous):

y=-5x because you substitute -5 for y and 1 for x so -5=a(1) and a=-5

OpenStudy (camerondoherty):

So \[\frac{ -5 }{ 1 }\] Is the slope, or -5 So, what you would put next to the x is -5

OpenStudy (camerondoherty):

So the equation would look like: \[y=-5x\]